Circular No. 7115

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

U SCORPII A. Niedzielski, Center for Astronomy, Torun University; and T. To-

mov and U. Munari, Padova and Asiago Astronomical Observatories, write: “We obtained spectra (range 435–495 nm; 0.1 nm/pixel) of U Sco (cf. IAUC 7113) around Feb. 26.14 UT with the Asiago 1.22-m telescope + Boller & Chivens CCD spectrograph). The spectrum of U Sco appears completely dominated by wide emission blends in which identification is made uncertain by the extreme broadness of the features (FWZI ∼ 8600–11 000 km/s). Probably present are N III, C III, C IV, and hydrogen emission lines. The presence of He I and He II lines is suspected, too. Monitoring continues.”

GK PERSEI This 1901 nova has been having a small outburst, the last having been

in 1996 (IAUC 6325). Visual magnitude estimates: Feb. 11.818 UT, 13.0 (P. Schmeer, Bischmisheim, Germany); 12.803, 13.0 (Schmeer); 14.87, 13.0 (Schmeer); 16.869, 13.0 (Schmeer); 24.796, 11.9 (Schmeer); 26.845, 11.8 (Schmeer); 27.758, 11.7 (K. Hornoch, Lelekovice, Czech Republic).

COMET C/1998 M5 (LINEAR)

Ephemeris extension to IAUC 7080 (elements MPC 32866):

1999 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Mar. 3 19^{h}42.^{m}8 +77^{o}29.^{′}6 1.537 1.808 88.^{o}6 33.^{o}2  9.5
5 19 45.1 +79 26.0 1.533 1.816 89.3 33.1  9.5
7 19 47.6 +81 24.4 1.529 1.823 90.0 33.0  9.5
9 19 50.7 +83 24.5 1.528 1.831 90.6 32.8  9.5
11 19 55  +85 26.2 1.528 1.839 91.2 32.7  9.6
13 20 04  +87 29.0 1.529 1.847 91.6 32.5  9.6
15 21 12  +89 31.4 1.532 1.856 92.0 32.4  9.6
17 7 29  +88 21.9 1.537 1.865 92.3 32.2  9.6
19 7 43  +86 17.6 1.544 1.874 92.5 32.0  9.7
21 7 49.0 +84 13.1 1.553 1.884 92.7 31.9  9.7
23 7 52.5 +82 09.1 1.563 1.893 92.7 31.7  9.7
25 7 55.3 +80 05.7 1.576 1.903 92.7 31.6  9.8
27 7 57.8 +78 03.4 1.590 1.914 92.5 31.4  9.8
29 8 00.1 +76 02.4 1.606 1.924 92.3 31.2  9.9
31 8 02.3 +74 03.0 1.623 1.935 92.0 31.0  9.9
Apr. 2 8 04.3 +72 05.5 1.642 1.947 91.6 30.9 10.0
1999 February 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green